A train leaves New York for Boston, 200 miles away, at 4:00 P.M. and averages 60 mph. Another train leaves Boston for New York on an adjacent set of tracks at 3:00 P.M. and averages 50 mph. At what time will the trains meet? (Round to the nearest minute.)
Answer by drk(1908)

This is an RTD problem. Here is a table based on the given information:
trains . . . . . . . . rate . . . . . . . . . .time . . . . . . . . . .distance
to boston . . . . . .60 . . . . . . . . . . .t . . . . . . . . . . . . . . .60t = 200
to NY . . . . . . . . .50 . . . . . . . . . . .t+1 . . . . . . . . . . . . .50t + 50 = 200
Looking at column 3, we want to know when they meet, so the sum of the distances = 200, or
60t + 50t + 50 = 200
110t + 50 = 200
110t = 150
t = 1.181818 hours.
which is ~ 1 hrs 22 min.
or about 4:22 PM